The electric held in a region is given by E =20i +30j in certain units. As one goes a small distance dl in the direction of
n vector =30i-20j .The potential will

(a) increase

(b) decrease

(c) remain constant

(d) change sign

Choose the correct option .More than one option could be correct
1
Expert's answer
2020-09-24T11:09:17-0400
"\Delta V=-\vec{E}\cdot\Delta \vec{s}=-(20,30)\begin{pmatrix}
 30 \\
 -20 
\end{pmatrix}\\=-(600-600)=0 \ V"

Answer: (c) remain constant